Find the vertex, the line of symmetry, the maximum or minimum value of the quadratic function and graph the function.
f(x)=12x^2-12x+4

Answer by nerdybill(7384)

Find the vertex, the line of symmetry, the maximum or minimum value of the quadratic function and graph the function.
f(x)=12x^2-12x+4
.
Line of symmetry:
x = -b/(2a)
x = -(-12)/(2*12)
x = 12/24
x = 1/2 (line of symmetry)
.
Vertex:
f(1/2) = 12(1/2)^2 - 12(1/2) + 4
f(1/2) = 12(1/4) - 6 + 4
f(1/2) = 3 - 6 + 4
f(1/2) = 1
vertex is at (1/2, 1)
.
Since the coefficient associated with the x^2 term is positive (think happy face) the vertex is a MINIMUM